princeevee:

Simplify. −2i√-12 Enter your answer, in simplest radical form,

princeevee:

@Mehek

princeevee:

@Ultrilliam

Mehek:

-2i√12 √12 = √4 * √3 √4 = 2 √12 = 2√3 -2i*2√3 What's -2 * 2 = ?

princeevee:

4

Mehek:

-4 -4i√3

Vocaloid:

@mehek I think you missed the negative sign inside the radical

Mehek:

Oh yeah I did